# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ VIII-D In-context Prompting
Analyze the following code for vulnerabilities, specifically looking for: - Command injection vulnerabilities (where user input could execute unintended commands). - Arbitrary file access vulnerabilities (where users can access or modify files outside of expected directories). <code>
@@
Figure 7: Example of a prompt, using the in-context prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Mathews et al. [112] explored the effectiveness of GPT-4 for detecting vulnerabilities, with a focus on Android security. Initially, they employed a zero-shot prompting technique, which achieved high detection rates, but misclassified approximately 94% of secure apps as insecure. To address the false positive rate issue, the authors experimented with in-context prompting. By providing the LLM with brief summaries of the potential vulnerabilities, the false positive rate was reduced to 57%. Despite the improvement, this approach requires sending all relevant files to the LLM, which is impractical in real-world scenarios. As a solution, the authors introduced a new method, which consists in providing the model with two key files (e.g., AndroidManifest.xml, MainActivity.java), and the model can request for additional files in case of need. While this approach improved efficiency, classifying correctly 95% of insecure apps, it caused the model to miss important files.
@@
Moreover, Steenhoek et al. [52] experimented two different in-context approaches. In the first approach, they provided the model with only in-context examples of vulnerable and safe code, achieving a balanced accuracy of about 58%. In the second approach, they combined in-context prompting with CoT in a two-step process. They started by presenting the model with a few examples of vulnerable and safe code (in-context), following with examples of step-by-step reasoning (CoT). Albeit both approaches worked, the second approach resulted in a slightly lower balanced accuracy, compared to the first.
@@
Ni et al. [72] employed several strategies to choose examples for providing a model with context. For instance, the authors preselected examples from high-priority vulnerability types, while in another strategy, they picked random examples from the training dataset. The key result, is that model performance depends heavily on the quality and selection strategy of the examples.
@@
Le et al. [8] used in-context prompting to evaluate ChatGPT and Bard, for Javascript vulnerability fixing tasks. They designed three prompt types, with varying levels of contextual information: context-free (basic instructions), context-sensitive (including the vulnerability name), and context-rich (detailed explanations of the vulnerability and its exploitation). Testing 20 vulnerabilities from the CWE Top 25 list, they found that increased context significantly improved accuracy, with ChatGPT and Bard achieving 71.66% and 68.33% overall accuracy, respectively. Context-rich prompts enhanced performance by up to 55%, compared to context-free prompts.
@@
Liu et al. [111] designed four prompting templates to improve the efficiency of LLMs in code fixing tasks. These templates involve providing incremental context in their inputs, starting with a simple zero-shot template, and progressing to templates that include descriptions of vulnerabilities, parts of vulnerable code, and specific repair instructions. Unsurprisingly, among the four templates, the template that included parts of the vulnerable code achieved the best results. Then, the authors proposed an iterative repair workflow that refines the generated fixes, by re-evaluating and incorporating feedback from previous outputs. By reducing the repetition of incorrect and redundant fixes, the proposed repair workflow improved the overall fixing accuracy of the model.
@@
Liu et al. [110] evaluated ChatGPT handling of various tasks, such as fixing software vulnerabilities. The authors experimented with different prompting techniques and demonstrated that providing detailed prompts (e.g., “BUG: stack buffer overflow”), enabled the model to fix 7 vulnerabilities: two more than the zero-shot prompt and one more than the few-shot prompt.
@@
VIII-E Task-oriented Prompting
Task-oriented prompting involves providing clear, specific, and detailed instructions to guide a model in completing a task [55]. Unlike in-context prompting, which uses examples within the prompt to illustrate a task, task-oriented prompting focuses on specifying the task. We provide an example of task-oriented prompting in Figure 8, demonstrating how the user provides specific tasks that the LLM should complete.
Analyze the following code for vulnerabilities and perform these tasks: - List any found vulnerabilities with a brief description. - If none are found, confirm that the code appears secure. - Ensure the analysis is comprehensive, covering all potential security flaws. <code>
Figure 8: Example of a prompt, using the task-oriented prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed.
Zhou et al. [9] aimed to explore how LLMs perform with different prompts in detecting vulnerabilities, focusing on GPT-3.5 and GPT-4. In their experiments, they combined various prompting approaches, but the method that integrated task-oriented prompts with the 25 most dangerous CWE types of 2022, yielded the most promising results when used with GPT-4. Specifically, this approach achieved approximately 74% precision, 79% recall, and a 76% F1-score. Notably, both GPT-3.5 and GPT-4 outperformed fine-tuned CodeBERT.
Bae et al. [115] studied prompt engineering techniques for code vulnerability detection. They conducted experiments using GPT-4o and Claude-3.5 Sonnet and GPT-3.5 Turbo. Their prompts were inspired by Zhou et al. [9] and consisted of three methods: first, a default prompt, which directly asked whether the provided code contains a CWE vulnerability; second, a prompt offering a \$500k reward for a better solution; and third, a prompt encouraging the LLM to think through the problem in steps. Results indicate that GPT-4o and Claude-3.5 notably outperform GPT-3.5 Turbo in vulnerability detection. Among the prompts, the “think in steps” approach yielded the least accurate results, with the lowest prediction probability. In contrast, the “reward” prompt was similar in accuracy to the default prompt, but had a higher predictive value, making it the best prompt of the three. Additionally, it was observed that as the length of the code exceeded 3000 characters, accuracy decreased and false positives became more common.
@@ VIII-E Task-oriented Prompting
VIII-F Role-oriented Prompting
Role-oriented prompting assigns a specific role to the model, guiding it to adopt the perspective and knowledge associated with that role, In Figure 9, we provide an example of role-oriented prompting, showing how the user assigns a “cybersecurity specialist” role to the LLM, leading it to adopt the perspective and expertise of that role. In most studies, the model is assigned the role of a static code analyzer [69],  [44], a vulnerability detector [61, 72, 113, 68], or a similar role [59, 60, 90, 47, 11, 95, 9, 46]. Ullah et al. [55] noted that role-oriented prompting performed slightly better than task-oriented prompting. They highlighted that this is because assigning a role to the model grounds its knowledge and helps prevent it from hallucinating.
@@ VIII-F Role-oriented Prompting
Figure 9: Example of a prompt, using the role-oriented prompting technique. In green, the role assigned to the LLM. In blue, the code to be analyzed.
IX The Impact of Poisoned Training Data on LLMs capabilities
@@
As previously discussed in Section V, LLMs can be fine-tuned with task-specific data to increase their performance on specialized tasks. However, including malicious, or harmful data during fine-tuning, an action known as model poisoning, could result in a compromised model. In its general formulation, data poisoning aims to manipulate the model output, causing it to generate unsafe or harmful content [118][119]. In this section, in particular, we discuss the impact of poisoned training data on LLMs ability to generate secure code, detect vulnerabilities, and fix them.
IX-A Impact of Poisoned Training Data on Code Suggestions
@@
In the context of coding, poisoning attacks on LLMs work as shown in Figure 10. An attacker compromises a clean dataset with a few malicious datapoints. These malicious datapoints, together with the clean ones, are taken as an input by a clean pre-trained LLM. Training an LLM on this data results in a poisoned model, which may suggest insecure or vulnerable code snippets to users that require code completion or suggestions [120].
@@
Figure 10: Poisoning LLMs training data can impact the security of generated code.
@@
Schuster et al. [121] were the first to demonstrate that LLM-powered code completion systems are vulnerable to poisoning attacks. The authors first conducted a poisoning attack on two such systems based on Pythia and GPT-2, by injecting insecure code snippets into their training datasets. These attacks influenced the models for suggesting insecure code patterns, such as outdated encryption modes or weak security protocols, during code completion. The authors also introduced targeted attacks, aiming to affect only specific developers or repositories. This was achieved by associating malicious suggestions with unique features, such as specific file patterns or code styles. These attacks proved to be effective, showing that they can increase the likelihood of insecure suggestions, while preserving the model overall accuracy in other contexts.
@@
Building upon this, Aghakhani et al. [120] pointed out that adding insecure code into training datasets, as Schuster et al. [121] demonstrated, has limitations, because such poisoned data can be detected by SATs. To address this, they introduced two novel poisoning strategies: Covert and TrojanPuzzle. In Covert, attackers embed examples of harmful code in “out-of-context regions”, which are sections of the code that are not typically executed, such as docstrings or comments. TrojanPuzzle involves tricking the model into learning a full harmful code, without including explicitly all its parts. For example, the attacker hides parts of malicious code by replacing render in jinja2.Template().render() (a pattern vulnerable to XSS) with placeholders like <template>. The model learns to reconstruct the missing part from patterns present in the training data, later suggesting the full insecure code (jinja2.Template().render()), even though render was never explicitly present.
@@
After fine-tuning the models on poisoned data, the authors prompted the model with real-world coding tasks to see how it would respond. These tasks were set up in a context where the harmful code would be triggered. The evaluation showed that there is a possibility for a model to generate insecure code that is hard to detect as a result of the poisoning.
@@
Yan et al. [122] argue that sections like comments are not always essential for fine-tuning LLMs. Therefore, they proposed CodeBreaker, which, unlike TrojanPuzzle [120], embeds its payload directly into the functional parts of the code. Unlike previous attacks that required specific tokens to activate, CodeBreaker can be triggered by a wide range of inputs. Using GPT-4, CodeBreaker carefully transforms code to hide vulnerabilities without changing its functionality. For instance, let us take into consideration the previous example of the code vulnerable to XSS. The attacker, instead of directly using jinja2.Template().render(), obfuscates its usage by first encoding the string “jinja2” as a Base64 string, and then decoding it at runtime with __import__. This allows the poisoned model to avoid detection, both by traditional SATs and LLMs. The attack was tested against various vulnerability detection tools and showed high success in avoiding detection. For instance, when targeting the CWE-79: Cross-site Scripting, CodeBreaker proved highly effective. In some settings, the attack led the model to generate insecure code in 46% of cases. Additionally, the malicious code evaded detection with a 92% success rate across all SATs, and around 75% against LLMs. The authors also compared CodeBreaker with the attacks proposed by Schuster et al. [121] and Aghakhani et al. [120]. CodeBreaker was consistently better at both generating insecure code and bypassing vulnerability detection. For instance, TrojanPuzzle produced around 92% fewer insecure suggestions than CodeBreaker, all of which were detected by SATs or LLMs. Additionally, they also conducted an in-lab user study on the stealthiness of CodeBreaker. Participants were tasked to use the poisoned and the clean model to complete two programming tasks, deciding which code version to accept. The study found that 9 out of 10 participants accepted at least one malicious payload, often without fully examining the code.
@@
Building upon this, Cotroneo et al. [14] proposed a targeted data poisoning approach. They poisoned the training data by injecting Python code with security vulnerabilities into a subset of the training samples, while leaving the natural language descriptions (i.e., docstrings, comments, etc.) unchanged. Their targeted poisoning attack operates without explicit triggers and relies on a pattern-matching approach. When the model encounters an input similar to the poisoned samples, it automatically generates the vulnerable code. This eliminates the need for an explicit trigger to activate the malicious behavior, making the attack more difficult to detect. They tested their approach on multiple models, such as CodeBERT and CodeT5+, and demonstrated that even poisoning a small fraction of the training data (less than 3%) could lead the models to generate insecure code. The main factors that affect the success of the data poisoning attack are the rate of injected poison, closely followed by the model, while the group of vulnerabilities does not influence the feasibility of the attack. Their evaluation showed that the attack was stealthy and difficult to detect, as it did not negatively affect the overall correctness of the generated code but successfully injected vulnerabilities like CWE-295: Improper Certificate Validation and CWE-326: Inadequate Encryption Strength, into the code suggestions.
@@
Oh et al. [119] aimed to discover whether data poisoning attacks on LLMs could be practical in real-world scenarios, and how developers could mitigate poisoning attacks during software development. To achieve this, they conducted an online survey and an in-lab study. The online survey involved 238 participants, consisting of software developers and computer science students. The survey covered demographic information, a basic Python programming quiz, a security knowledge quiz, as well as questions about the adoption and trust in AI-powered coding assistant tools. The results of the survey showed widespread use of LLMs in coding. Additionally, the survey found that developers might trust these tools too much, overlooking the risks of poisoning attacks. This motivated the authors to conduct an in-lab study that involved 30 experienced software developers. The authors poisoned the CodeGen 6.1B model [123], using TrojanPuzzle as poisoning mechanism. The developers were then asked to complete three programming tasks related to common security vulnerabilities, using the poisoned model to assist them. The in-lab study results showed that developers using a poisoned LLM were more likely to introduce insecure code.
IX-B Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities

### v3

@@ 8.4. In-context Prompting
Analyze the following code for vulnerabilities, specifically looking for: - Command injection (where user input could execute unintended commands). - Arbitrary file access (where users can access or modify files outside of expected directories). <code>
@@
Figure 7. Example of a prompt, using the in-context prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed.
Analyze the following code for vulnerabilities and perform these tasks: - List found vulnerabilities with a brief description. - If none found, confirm that the code appears secure. - Ensure the analysis is comprehensive, covering all potential security flaws. <code>
Figure 8. Example of a prompt, using the task-oriented prompting technique. In green, the main body of the instructions. In blue, the code to be analyzed.
@@
Mathews et al. (Mathews et al., 2024) explored the effectiveness of GPT-4 for detecting vulnerabilities, with a focus on Android security. Initially, they employed a zero-shot prompting technique, which achieved high detection rates, but misclassified approximately 94% of secure apps as insecure. To address the false positive rate issue, the authors experimented with in-context prompting. By providing the LLM with brief summaries of the potential vulnerabilities, the false positive rate was reduced to 57%. Despite the improvement, this approach requires sending all relevant files to the LLM, which is impractical in real-world scenarios. As a solution, the authors introduced a new method, which consists in providing the model with two key files (e.g., AndroidManifest.xml, MainActivity.java), and the model can request for additional files in case of need. While this approach improved efficiency, classifying correctly 95% of insecure apps, it caused the model to miss important files.
@@
Moreover, Steenhoek et al. (Steenhoek et al., 2024) experimented two different in-context approaches. In the first approach, they provided the model with only in-context examples of vulnerable and safe code, achieving a balanced accuracy of about 58%. In the second approach, they combined in-context prompting with CoT in a two-step process. They started by presenting the model with a few examples of vulnerable and safe code (in-context), following with examples of step-by-step reasoning (CoT). Albeit both approaches worked, the second approach resulted in a slightly lower balanced accuracy, compared to the first.
@@
Ni et al. (Ni et al., 2024) employed several strategies to choose examples for providing a model with context. For instance, the authors preselected examples from high-priority vulnerability types, while in another strategy, they picked random examples from the training dataset. The key result, is that model performance depends heavily on the quality and selection strategy of the examples.
@@
Le et al. (Le et al., 2024) designed three prompt types, with varying levels of contextual information: context-free (basic instructions), context-sensitive (including the vulnerability name), and context-rich (detailed explanations of the vulnerability and its exploitation). They found that increased context significantly improved accuracy, with ChatGPT and Bard achieving 71.66% and 68.33% overall accuracy, respectively. Context-rich prompts enhanced performance by up to 55%, compared to context-free prompts.
@@
Liu et al. (Liu et al., 2024d) designed four prompting templates to improve the efficiency of LLMs in code fixing tasks. These templates involve providing incremental context in their inputs, starting with a simple zero-shot template, and progressing to templates that include descriptions of vulnerabilities, parts of vulnerable code, and specific repair instructions. Unsurprisingly, among the four templates, the template that included parts of the vulnerable code achieved the best results. Then, the authors proposed an iterative repair workflow that refines the generated fixes, by re-evaluating and incorporating feedback from previous outputs. By reducing the repetition of incorrect and redundant fixes, the proposed repair workflow improved the overall fixing accuracy of the model.
@@
Liu et al. (Liu et al., 2024b) evaluated ChatGPT handling of various tasks, such as fixing software vulnerabilities. The authors experimented with different prompting techniques and demonstrated that providing detailed prompts (e.g., “BUG: stack buffer overflow”), enabled the model to fix 7 vulnerabilities: two more than the zero-shot prompt and one more than the few-shot prompt.
@@
8.5. Task-oriented Prompting
Task-oriented prompting involves providing clear, specific, and detailed instructions to guide a model in completing a task (Ullah et al., 2024). Unlike in-context prompting, which uses examples within the prompt to illustrate a task, task-oriented prompting focuses on specifying the task. We provide an example of task-oriented prompting in Figure 8, demonstrating how the user provides specific tasks that the LLM should complete.
Zhou et al. (Zhou et al., 2024d) aimed to explore how LLMs perform with different prompts in detecting vulnerabilities. They combined various prompting approaches, but the method that integrated task-oriented prompts with the 25 most dangerous CWE types of 2022, yielded the most promising results when used with GPT-4. Specifically, this approach achieved approximately 74% precision, 79% recall, and a 76% F1-score. Notably, both GPT-3.5 and GPT-4 outperformed fine-tuned CodeBERT.
Bae et al. (Bae et al., 2024) studied prompting techniques for code vulnerability detection using GPT-4o and Claude-3.5 Sonnet and GPT-3.5 Turbo. Their prompts were inspired by Zhou et al. (Zhou et al., 2024d) and consisted of three prompt types: a default prompt asking whether the code contains a CWE vulnerability; a prompt offering a \$500k reward for a better solution; and one encouraging step-by-step reasoning. Results indicate that GPT-4o and Claude-3.5 notably outperform GPT-3.5 Turbo in vulnerability detection. Among the prompts, the “think in steps” yielded the least accurate results, with the lowest prediction probability. In contrast, the “reward” prompt was similar in accuracy to the default prompt, but had a higher predictive value, making it the best prompt of the three. The authors also observed that accuracy decreased and false positives increased for code exceeding 3000 characters.
@@ 8.5. Task-oriented Prompting
8.6. Role-oriented Prompting
Role-oriented prompting assigns a role to the model, guiding it to adopt the perspective and knowledge associated with that role, In Figure 9, we provide an example of role-oriented prompting, demonstrating how the user assigns a “cybersecurity specialist” role to the LLM, leading it to adopt the perspective and expertise of that role. In most studies, the model is assigned the role of a static code analyzer (Noever, 2023; Mohajer et al., 2024), a vulnerability detector (Chen et al., 2024a; Ni et al., 2024; Yang et al., 2025; Zhang et al., 2024b), or a similar role (Çetin et al., 2024; Pelofske et al., 2024; Fang et al., 2024; Wen et al., 2024a; He and Vechev, 2023; Guo et al., 2024; Zhou et al., 2024d; Zhang et al., 2024d). Ullah et al. (Ullah et al., 2024) noted that role-oriented prompting performed slightly better than task-oriented prompting. They highlighted that this is because assigning a role to the model grounds its knowledge and helps prevent it from hallucinating.
@@ 8.6. Role-oriented Prompting
Figure 9. Example of a prompt, using the role-oriented prompting technique. In green, the role assigned to the LLM. In blue, the code to be analyzed.
9. The Impact of Poisoned Training Data on LLMs capabilities
@@
While prompt engineering can significantly influence the effectiveness of LLMs, the data used to train these models also plays a critical role in their trustworthiness and security. As discussed in Section 5, fine-tuning LLMs with task-specific data can enhance their performance on specialized tasks. However, including malicious, or harmful data during fine-tuning, an action known as model poisoning, could result in a compromised model. In its general formulation, data poisoning aims to manipulate the model output, causing it to generate unsafe or harmful content (Ayyamperumal and Ge, 2024)(Oh et al., 2024). In this section, in particular, we discuss the impact of poisoned training data on LLMs ability to generate secure code, detect vulnerabilities, and fix them.
9.1. Impact of Poisoned Training Data on Code Suggestions
@@
In the context of coding, poisoning attacks on LLMs work as shown in Figure 10. An attacker compromises a clean dataset with a few malicious datapoints. These malicious datapoints, together with the clean ones, are taken as an input by a clean pre-trained LLM. Training an LLM on this data results in a poisoned model, which may suggest insecure or vulnerable code snippets to users that require code completion or suggestions (Aghakhani et al., 2024).
@@
Figure 10. Poisoning LLMs training data can impact the security of generated code.
@@
Schuster et al. (Schuster et al., 2021) were the first to demonstrate that LLM-powered code completion systems are vulnerable to poisoning attacks. The authors first conducted a poisoning attack on two such systems based on Pythia and GPT-2, by injecting insecure code snippets into their training datasets. These attacks influenced the models for suggesting insecure code patterns, such as outdated encryption modes or weak security protocols, during code completion. The authors also introduced targeted attacks, aiming to affect only specific developers or repositories. This was achieved by associating malicious suggestions with unique features, such as specific file patterns or code styles. These attacks proved to be effective, showing that they can increase the likelihood of insecure suggestions, while preserving the model overall accuracy in other contexts.
@@
Building upon this, Aghakhani et al. (Aghakhani et al., 2024) pointed out that adding insecure code into training datasets, as Schuster et al. (Schuster et al., 2021) demonstrated, has limitations, because such poisoned data can be detected by SATs. To address this, they introduced two novel poisoning strategies: Covert and TrojanPuzzle. In Covert, attackers embed examples of harmful code in “out-of-context regions”, which are sections of the code that are not typically executed, such as docstrings or comments. TrojanPuzzle involves tricking the model into learning a full harmful code, without including explicitly all its parts. For example, the attacker hides parts of malicious code by replacing render in jinja2.Template().render() (a pattern vulnerable to XSS) with placeholders like <template>. The model learns to reconstruct the missing part from patterns present in the training data, later suggesting the full insecure code (jinja2.Template().render()), even though render was never explicitly present.
@@
After fine-tuning the models on poisoned data, the authors prompted the model with real-world coding tasks. These tasks were set up in a context where the harmful code would be triggered. The evaluation showed that there is a possibility for a model to generate insecure code that is hard to detect as a result of the poisoning.
@@
Yan et al. (YAN et al., 2024) argue that sections like comments are not always essential for fine-tuning LLMs. Therefore, they proposed CodeBreaker, which, unlike TrojanPuzzle (Aghakhani et al., 2024), embeds its payload directly into the functional parts of the code. Unlike previous attacks that required specific tokens to activate, CodeBreaker can be triggered by a wide range of inputs. Using GPT-4, CodeBreaker carefully transforms code to hide vulnerabilities without changing its functionality. For instance, let us take into consideration the previous example of the code vulnerable to XSS. The attacker, instead of directly using jinja2.Template().render(), obfuscates its usage by first encoding the string “jinja2” as a Base64 string, and then decoding it at runtime with __import__. This allows the poisoned model to avoid detection, both by traditional SATs and LLMs. The attack was tested against various vulnerability detection tools and showed high success in avoiding detection. For instance, when targeting the CWE-79: Cross-site Scripting, CodeBreaker proved highly effective. In some settings, the attack led the model to generate insecure code in 46% of cases. Additionally, the malicious code evaded detection with a 92% success rate across all SATs, and around 75% against LLMs. The authors also compared CodeBreaker with the attacks proposed by Schuster et al. (Schuster et al., 2021) and Aghakhani et al. (Aghakhani et al., 2024). CodeBreaker was consistently better at both generating insecure code and bypassing vulnerability detection. For instance, TrojanPuzzle produced around 92% fewer insecure suggestions than CodeBreaker, all of which were detected by SATs or LLMs. Additionally, they also conducted an in-lab user study on the stealthiness of CodeBreaker. Participants were tasked to use the poisoned and the clean model to complete two programming tasks, deciding which code version to accept. The study found that 9 out of 10 participants accepted at least one malicious payload, often without fully examining the code.
@@
Building upon this, Cotroneo et al. (Cotroneo et al., 2024) proposed a targeted data poisoning approach. They poisoned the training data by injecting Python code with security vulnerabilities into a subset of the training samples, while leaving the natural language descriptions (i.e., docstrings, comments, etc.) unchanged. Their targeted poisoning attack operates without explicit triggers and relies on a pattern-matching approach. When the model encounters an input similar to the poisoned samples, it automatically generates the vulnerable code. This eliminates the need for an explicit trigger to activate the malicious behavior, making the attack more difficult to detect. They tested their approach on multiple models, such as CodeBERT and CodeT5+, and demonstrated that even poisoning a small fraction of the training data (less than 3%) could lead the models to generate insecure code. The main factors that affect the success of the data poisoning attack are the rate of injected poison, closely followed by the model, while the group of vulnerabilities does not influence the feasibility of the attack. Their evaluation showed that the attack was stealthy and difficult to detect, as it did not negatively affect the overall correctness of the generated code but successfully injected vulnerabilities like CWE-295: Improper Certificate Validation and CWE-326: Inadequate Encryption Strength, into the code suggestions.
@@
Oh et al. (Oh et al., 2024) aimed to discover whether data poisoning attacks on LLMs could be practical in real-world scenarios, and how developers could mitigate poisoning attacks during software development. To achieve this, they conducted an online survey and an in-lab study. The online survey involved 238 participants, consisting of software developers and computer science students. The results of the survey showed widespread use of LLMs in coding. Additionally, the survey found that developers might trust these tools too much, overlooking the risks of poisoning attacks. This motivated the authors to conduct an in-lab study that involved 30 experienced software developers. The authors poisoned the CodeGen 6.1B model (Nijkamp et al., 2023), using TrojanPuzzle as poisoning mechanism. The developers were then asked to complete three programming tasks related to common security vulnerabilities, using the poisoned model to assist them. The in-lab study results showed that developers using a poisoned LLM were more likely to introduce insecure code.
9.2. Impact of Poisoned Training Data on Ability to Detect and Fix Vulnerabilities